OpenStudy (anonymous):

The volume of 2 similar solids are 27 ft3 and 216 ft3. The surface area of the larger solid is 300 ft2. What is the surface area of the smaller solid? Round your answer to the nearest tenths place.

OpenStudy (anonymous):

Volume ratio = 216/27, and is the cube of the linear scale factor. So, each side of the larger is 6/3 = 2 times the smaller. So, the area of the larger is 2^2 = 4 times the smaller.

OpenStudy (anonymous):

Thanks britney spears lol but my answer choices are so different: a. 13.7 ft2 c. 37.5 ft2 b. 1,200.0 ft2 d. 75.0 ft2

OpenStudy (anonymous):

i tried doing what you did though

OpenStudy (anonymous):

The cube root of 216 = 6; the cube root of 27 = 3 So the ratio of the linear dimensions is 6:3 or 2:1 Now the area varies as the square of the linear dimensions, so here the ratio is (2^2):(1^2), or 4:1 So the area of the smaller solid is 75 sq ft

OpenStudy (anonymous):

I made a mistake on my first answer, sorry about that!!

OpenStudy (anonymous):

ohh! okay thanks! :) its okay lol you made it up. thanks britney!

OpenStudy (anonymous):

You're welcome!! Xo!